Calculate the ratio of ammonium chloride to ammonia that is required to make a buffer solution with a pH of 9.00. The Ka for ammonium ion is 5.6 x 10-10.

1
Expert's answer
2022-05-17T10:11:04-0400

Solution:

The pH of the buffer solution can be calculated from the Henderson-Hasselbalch equation:



where the base is ammonia (NH3) and the acid is ammonium chloride (NH4Cl).


You can rearrange this equation to solve for the ratio of ammonia to ammonium chloride:

[NH3] / [NH4Cl] = 10(pH − pKa)


pKa = −log(Ka) = −log(5.6×10−10) = 9.25

pH = 9.00

Therefore,

[NH3] / [NH4Cl] = 10(9.00 − 9.25) = 10−0.25 = 0.5623

[NH3] / [NH4Cl] = 0.5623


Calculate the ratio of ammonium chloride to ammonia:

[NH4Cl] / [NH3] = (0.5623)−1 = 1.78

[NH4Cl] / [NH3] = 1.78


Answer: The ratio of ammonium chloride to ammonia is 1.78